How to slove this Quadratic equation by completing the square: x^2+4x-8=0
Answer by rapaljer(4671)

x^2 + 4x +_____ = 8 + _____

Take HALF of the x term, and square the result. In this case, half of 4 is 2, and 2 squared is 4. So add +4 to each side:

x^2 + 4x + 4 = 8 + 4

Note Perfect Square Trinomial on the left side:
(x+2)^2 = 12

Square root each side:
(x+2) = 0+-sqrt(12)
x+2 = 0 +- sqrt(4)*sqrt(3)

Subtract 2 from each side:
x= -2 +- 2*sqrt(3)

R^2